Can you please help me with the question 2(3-x)=14 what is the value of x? I need a really good explanation please I need it as fast as possible and if possible right now.
Answer by mananth(16946)

2(3-x)=14

divide the equation by 2
2(3-x)/2 = 14/2
3-x = 7
subtract 3 from both sides
3-x-3 = 7-3
-x=4
divide by -1
x= -4